OpenStudy (anonymous):

help me solve: (a^2)^-5

hartnn (hartnn):

\(\huge (x^n)^m=x^{mn}\) does this help?

OpenStudy (anonymous):

so the answer is a^10?

hartnn (hartnn):

2*(-5)=-10 so its a^(-10) ok?

OpenStudy (anonymous):

so I don't flip the exponent and make it positive?

hartnn (hartnn):

nopes, not here. you flip the sign of exponent only when you take its reciprocal \(\huge \frac{1}{x^n}=x^{-n}\)

OpenStudy (anonymous):

OHHHHH, okay- that helps. ty!

hartnn (hartnn):

welcome ^_^